OpenStudy (anonymous):

Find the difference quotient for f(x)=12/x^2 Any help is greatly appreciated.

OpenStudy (anonymous):

\[f(x)=\frac{ 12 }{ x^2 } \]

OpenStudy (anonymous):

Im still struggling with this one I just dont get it please help!

OpenStudy (zzr0ck3r):

\[\frac{\frac{12}{(x+h)^2}-\frac{12}{x^2}}{h}\]

OpenStudy (anonymous):

:) that is exactly where I am stuck.....

OpenStudy (zzr0ck3r):

\[\frac{\frac{12x^2}{(x+h)^2*x^2}-\frac{12(x+h)^2}{(x+h)^2x^2}}{h}\]= \[\frac{12x^2-(x^2+2xh+h^2)}{h(x+h)^2x^2}\]= \[\frac{11x^2-2xh-h^2}{h(x+h)^2x^2}\]

OpenStudy (zzr0ck3r):

do you need to take the limit?

OpenStudy (zzr0ck3r):

woops sec I messed up

OpenStudy (anonymous):

interesting so you did multiplied side by side.... and I think I need the limit. Which I believe is the simplest form?

OpenStudy (zzr0ck3r):

\[\frac{12x^2-12(x^2+2xh+h^2)}{hx^2(x+h)^2}\]= \[\frac{12x^2-12x^2-24xh-12h^2}{hx^2(x+h)^2}\]= \[\frac{-24xh-12h^2}{hx^2(x+h)^2}\]= \[\frac{-24-12h}{x^2(x+h)^2}\]

OpenStudy (zzr0ck3r):

when I say limit, I mean when h goes to infinity. Is this for calculus or algebra?

OpenStudy (anonymous):

Alg

OpenStudy (zzr0ck3r):

I mean 0 not infinity

OpenStudy (zzr0ck3r):

ok well then you don't want the limit im guessing.

OpenStudy (zzr0ck3r):

this is the form you want.

OpenStudy (zzr0ck3r):

you are actually doing calculus you just don't know it yet:)

OpenStudy (anonymous):

God i can only imagine the rest....

OpenStudy (anonymous):

either way I chose this path so better get used to it :) one last question... Is this the simplest form I am looking at other exmaples and they're nice and neat

OpenStudy (anonymous):

Out of all this one was vicious

OpenStudy (zzr0ck3r):

\[\frac{-24x-12h}{x^2(x+h)^2}\]

OpenStudy (zzr0ck3r):

that's the final

OpenStudy (anonymous):

K K thank you so much I will try to sort the others by myself... :) looks like another all nighter. Thanks again

OpenStudy (zzr0ck3r):

yeah there is not much else you can do to it http://www.wolframalpha.com/input/?i=simplify++%28%E2%88%9224x%E2%88%9212h+%29%2F%28x%5E2*%28x%2Bh%29%5E2%29++++

OpenStudy (anonymous):

nice. Will check this site little bit more.